A population has a mean of 15 and a standard deviation of 3. What is the probability that a sample of 4 individuals will have a mean between 16.5 and 18? (2 marks)
Answer #1

Solution :

Given that,

mean = = 15

standard deviation = = 3

= / n = 3 / 4 = 1.5

= P[(16.5 - 15) /1.5 < ( - ) / < (18 - 15) /1.5 )]

= P(1 < Z < 2)

= P(Z < 2) - P(Z < 1)

= 0.9772 - 0.8413

= 0.1359

Probability = 0.1359